OpenStudy (anonymous):

If x-y=y-z=z-t=5, what is the value of x-t

OpenStudy (anonymous):

\[z = t+5 \] \[y = z+5 = t+10 \] \[x = y+5 = t+15 \] \[x-t = 15 \]

OpenStudy (anonymous):

Awesome , thanks bro!!!!!!